OpenStudy (ria23):

Given the table below, evaluate the following: x -3 -2 -1 0 1 2 3 4 5 f(x) 10 20 30 40 50 60 70 80 90 g(x) -1 -2 -3 -4 -5 -6 -7 -8 -9 a. (3f+2g)(1) b. (f o g)(-1)

jimthompson5910 (jim_thompson5910):

hint: (3f+2g)(1) = 3*f(1) + 2*g(1)

OpenStudy (ria23):

So... 3*10*1 + 2*-1*1, 30+(-2)?

jimthompson5910 (jim_thompson5910):

what is f(1) equal to?

jimthompson5910 (jim_thompson5910):

to find out, locate x = 1 and then look directly below it (to find the f(x) row)

OpenStudy (ria23):

50?

jimthompson5910 (jim_thompson5910):

since f(1) = 50, this means 3*f(1) = 3*50

jimthompson5910 (jim_thompson5910):

same logic happens with g(1)

OpenStudy (ria23):

So 3*(-5)(1) for g?

jimthompson5910 (jim_thompson5910):

g(1) represents the output when you plug in x = 1 into the g(x) function

jimthompson5910 (jim_thompson5910):

it's not g times 1

jimthompson5910 (jim_thompson5910):

g(1) is all together one number

OpenStudy (ria23):

Oh! So it would just be 3*-5...

OpenStudy (ria23):

2*-5**

OpenStudy (ria23):

My bad.

jimthompson5910 (jim_thompson5910):

2*(-5)

jimthompson5910 (jim_thompson5910):

so... (3f+2g)(1) = 3*f(1) + 2*g(1) (3f+2g)(1) = 3*50 + 2*(-5) (3f+2g)(1) = ???

OpenStudy (ria23):

150+(-10) 140...

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (ria23):

^o^ Awesome! Thank yhu! c: How would I work B out?

jimthompson5910 (jim_thompson5910):

(f o g)(-1) = f( g(-1) ) so you first need to compute g(-1) that result will be plugged into f(x) to get your final answer

jimthompson5910 (jim_thompson5910):

g(-1) = ???

OpenStudy (ria23):

-3

jimthompson5910 (jim_thompson5910):

therefore, f( g(-1) ) = f( -3 )

jimthompson5910 (jim_thompson5910):

notice how I replaced g(-1) with -3 the two are equal/equivalent so it's valid

jimthompson5910 (jim_thompson5910):

now you just need to find f( -3 )

OpenStudy (ria23):

-3*30?

jimthompson5910 (jim_thompson5910):

locate -3 on the x row then look down to the f(x) row

jimthompson5910 (jim_thompson5910):

you're thinking that f(-3) is f times -3 which is wrong

OpenStudy (ria23):

Oh! it would be 10

jimthompson5910 (jim_thompson5910):

yes, f(-3) is the same as the number 10

jimthompson5910 (jim_thompson5910):

input: x = -3 output: f(-3) = 10 rule: unknown function

jimthompson5910 (jim_thompson5910):

so that means f( g(-1) ) = 10 and (f o g)(-1) = 10

OpenStudy (ria23):

And that's the answer for B? o.o Woah... That was actually kind of easy.. n_n

jimthompson5910 (jim_thompson5910):

yeah it's not too bad

OpenStudy (ria23):

Thank yhu so much for the help!